OpenStudy (anonymous):

a punch recipe that calls for 2 liters of club soda and three 750 milliliter bottles of sparkling apple cider. How much of these two ingredients does the recipe call for?

OpenStudy (anonymous):

5 liters 2 deciliters 5 centiliters 5 liters 2 centiliters 5 milliliters 4 liters 2 centiliters 5 milliliters 4 liters 2 deciliters 5 centiliters

OpenStudy (anonymous):

those are my answer i have to choose from

OpenStudy (anonymous):

2 liters + 3*750 milliliters 2 liters + 2250 milliliters Now convert the milliliters to liters.

OpenStudy (anonymous):

There are 1000 milliliters in 1 liter, 100 centiliters in a liter, and 10 deciliters in a liter. the recipe calls for two 750 milliliter bottles. 2x750=1500 milliliters, or 15 deciliters. Since you already need 1 liter of club soda, and there are 10 deciliters in a liter, you have 1 liter and 5 deciliters of sparkling apple cider+1 liter of club soda=2 liters and 5 deciliters. So the recipe calls for 2 liters and 5 deciliters.

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!!!!!!!!!!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

5 liters 2 deciliters 5 centiliters 5 liters 2 centiliters 5 milliliters 4 liters 2 centiliters 5 milliliters 4 liters 2 deciliters 5 centiliters which answer?

OpenStudy (anonymous):

it would be a

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

I have another qusetion...While Stacy warmed up a microwave meal for her lunch, she took a look at the nutrition facts label. Her meal has 11 grams of protein and 600 milligrams of sodium. How much more protein is in her lunch than sodium?

OpenStudy (anonymous):

I am fairly certain that the answer to the first question is d, not a. Let me recheck my math.

OpenStudy (anonymous):

it was d... but that is ok

OpenStudy (anonymous):

10 grams 4 decigrams 10 grams 4 centigrams 11 grams 4 decigrams 11 grams 4 centigrams these are the answers i have to choose from for my second question.

OpenStudy (anonymous):

I have another qusetion...While Stacy warmed up a microwave meal for her lunch, she took a look at the nutrition facts label. Her meal has 11 grams of protein and 600 milligrams of sodium. How much more protein is in her lunch than sodium?

OpenStudy (anonymous):

The first step to solving the problem would be to convert the milligrams to grams. 600 milligrams = ? grams?

OpenStudy (anonymous):

aren't there 1000 milligrams in grams?

OpenStudy (anonymous):

0.6 right?

OpenStudy (anonymous):

610 milligrams is 0.610 grams, and 14 grams is 14 000 milligrams. The easiest way to do this would be like this: 14 000 - 610 _______ 13 390 = 13 grams, 3 decigrams, 9 centigrams.

OpenStudy (anonymous):

it is 600 milligrams,

OpenStudy (anonymous):

would it be, 10 grams 4 centigrams

OpenStudy (anonymous):

Not quite, but you're close! 600 mg would not even equal a full gram. But it would equal 6 decigrams. How many more decigrams would you need to make a gram?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

so it would be 11grams and 4 decigrams

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

10 grams 4 decigrams

OpenStudy (anonymous):

Yes. But do you understand why you got that answer? If not, I can continue walking you through the steps to solve it.

OpenStudy (anonymous):

yea, i noticed what i did wrong with my math. i did for about that even though there is 11 grams i would still need two subtract. Thank you so much for your help!!!!!! :)

OpenStudy (anonymous):

No problem! :)

OpenStudy (mathstudent55):

The original question was: A punch recipe that calls for 2 liters of club soda and three 750 milliliter bottles of sparkling apple cider. How much of these two ingredients does the recipe call for? A. 5 liters 2 deciliters 5 centiliters B. 5 liters 2 centiliters 5 milliliters C. 4 liters 2 centiliters 5 milliliters D. 4 liters 2 deciliters 5 centiliters Solution: 2 L + 3 * 750 mL = 2 L + 2250 mL = 2000 mL + 2250 mL = 4250 mL = 4 liters + 2 deciliters + 5 centiliters The answer is D.

OpenStudy (mathstudent55):

The second question was: While Stacy warmed up a microwave meal for her lunch, she took a look at the nutrition facts label. Her meal has 11 grams of protein and 600 milligrams of sodium. How much more protein is in her lunch than sodium? A. 10 grams 4 decigrams B. 10 grams 4 centigrams C. 11 grams 4 decigrams D. 11 grams 4 centigrams Solution: 11 g - 600 mg = 11g - 0.6 mg = 10.4 g = 10 grams + 4 decigrams Answer is A.